The number of transitive closure exists in the relation R = {(0,1), (1,2), (2,2), (3,4), (5,3), (5,4)} where {1, 2, 3, 4, 5} ∈ A is__________.

1
Expert's answer
2022-03-10T02:55:55-0500

Solution


Given that "R = \left \{ (0,1), (1,2), (2,2), (3,4), (5,3), (5,4) \right \}"


We consider a relation on a set "A", let it be "R".


Then the connectivity relation on "R^{*}" will consist of the pairs of the form "(a, b)" , with this condition that there is the path length of at least one "(=1)" from "a" to "b".


We represent it mathematically as,


"R^{*}=R^{1}\cup R^{2}\cup R^{3}....\cup R^{n}"


Therefore, the answer for the given blank space is


"R = \left \{ 1, 2, 3, 4, 5 \right \}∈ A" is "\left \{ (0,1), (0,2), (1,2), (2,2), (3,4), (5,3), (5,4) \right \}"